THE STUDY OF INTERMOLECULAR INTERACTIONS IN LIQUID BUTYRIC ACID AND ITS SOLU-TIONS BY RAMAN SPECTRA

A. JUMABAEV, F. H. TUKHVATULLIN, \underline{U. N. TASHKENBAEV}, Z. MAMATOV and S. A. OS-MANOV, Samarkand State University, 703004 Samarkand, Univ. Blvd, 15, Uzbekistan (email: info@samarkand.uz).